### v2

# Log-Normal Widths of Waiting Time Distributions: A Fundamental Parameter

Abstract
Many astronomical phenomena, including Fast Radio Bursts and Soft Gamma Repeaters, consist of brief, separated, seemingly aperiodic events. The intervals between these events vary randomly, but there are epochs of greater activity, with shorter mean intervals, and of lesser activity, with longer mean intervals. This variability can be quantified by a single dimensionless parameter, the width of a log-normal fit to the distribution of waiting times between events. If the distribution of event strengths is a power law, as is often the case, this parameter is independent of the detection threshold, and is a robust measure of the intrinsic variability of the waiting times and of the underlying dynamics.
1 Introduction
Many episodic natural phenomena originate in complex and imperfectly understood physical processes e.g., repeating Fast Radio Bursts (FRB) (Zhang, 2023) and Soft Gamma Repeaters (SGR) (Hurley, 2011). In some, such as FRB, the responsible physical processes and their environment are not known. In others we may know which physical processes are responsible (magnetic reconnection in SGR), but lack sufficient understanding to calculate them quantitatively.
Burst activity may be described by the distribution of intervals (waiting times) between detected bursts. This reflects two distinct properties of the underlying process: the correlation (or lack thereof) between consecutive bursts (short-term memory) and slower variations in the mean level of activity (long-term memory). When the level of activity varies, the width of the distribution increases; SGR 1806--20 is an example (Table 1). The distributions of waiting times in repeating FRB (Katz, 2019; Aggarwal et al., 2022; Hewitt et al., 2022; Li et al., 2021; Niu et al., 2022; Zhang et al., 2022) and in SGR burst storms (Hurley et al., 1994; Younes et al., 2020) have been well fit by log-normal functions. This paper describes the application of log-normal fits to waiting time distributions, and suggests their dimensionless standard deviation $\sigma$ as a robust and fundamental quantitative metric. Alternative statistical methods have recently been used to describe FRB waiting times by Du et al. (2023); Wang et al. (2023). Their conclusions about memory in burst recurrence intervals (“short waiting times tend to be followed by short ones, and long by long”) are qualitatively consistent with the picture of varying levels of activity presented here.
2 Bursting Phenomena
Bursting phenomena may be described by their distribution of strengths (energy, flux, fluence, electromagnetic field) and by their temporal distribution. Distributions of strengths are power laws (Younes et al., 2020; Zhang et al., 2022) when there is no characteristic value of strength (earthquake magnitude, flare or burst energy, etc.) or scale over a wide range (Kolmogorov, 1941a, b; Katz, 1986). Any deviation from a power law would define a characteristic strength, the value of strength where its distribution deviates from a straight line on a log-log plot, contradicting the assumption that there is no characteristic strength. In some FRB a break (a change in the exponent of the fitted energy distribution $f(E)$) is observed (Zhang et al., 2022), defining a characteristic value of the burst energy $E$ or fluence where the exponent changes, and some SGR have shown extreme outliers, also not described by a power law distribution (Katz, 2021); power laws are widespread but not universal. As a result, many but not all log-normal widths are independent of detection sensitivity.
Most episodically outbursting phenomena display periods of greater and lesser activity with shorter and longer mean waiting times, respectively. If the activity level is uniform the process is, by definition, Poissonian and the width of the waiting time distribution $\sigma=0.723$, while if there are periods of greater and lesser activity there are excesses of short and long waiting times, leading to a broader distribution of waiting times. The width $\sigma$ of this distribution thus measures the variability of the activity. In one limit outbursts are regularly periodic and $\sigma=0$, while in the other limit brief periods of frequent activity are separated by long periods of quietude and $\sigma\gg 1$; shot noise, with random events occurring at a constant mean rate, is intermediate.
3 Log-Normal Fits
Log-normal fitting functions are widely used because with only three parameters, a midpoint, a maximum and a width, they provide good fits to a broad range of single-peaked distributions, wide as well as narrow (, 1994; Limpert, Stahel & Abbt, 2001). These fits may be entirely phenomenological, without basis in a causal model. Most uses of log-normal fits describe the distribution of some parameter of the individual events, but here we are concerned with the waiting times between them.
The logarithm of the product of a series of multiplications of a random variable is the sum of the logarithms of the variables, and executes a random walk. If there are many factors in the product the central limit theorem applies to the sum of the logarithms, and the result approaches a log-normal distribution (Shockley, 1957; Montroll & Shlesinger, 1982). The rate of convergence depends on the distributions of these logarithms, and for pathological distributions the sum may not converge at all.
A log-normal distribution of a variable is therefore a natural consequence if it is the result of a series of independent stochastic multiplicative steps. The observed $\sigma$ of a waiting time distribution would be $\propto N^{-1/2}$, where $N$ is the number of independent steps, in series, weighted by their frequency: if there is a necessary step that occurs at a low rate, the observed waiting time distribution would reflect the distribution of that infrequent step, while a step that occurs at a high rate would have little effect on $\sigma$. A log-normal distribution will be a good fit if there are several required steps, each with a comparable rate, making the central limit theorem applicable to the sum of their logarithms.
Power-law distributions are not fit by log-normal functions because they do not have a peak. All power-law distributions of data must have at least one cutoff, either a threshold for detection or an intrinsic characteristic scale, in order that the total number of events and the total energy (or an analogous quantity) be finite. The distribution peaks at the cutoff because it marks the transition between an increasing function that would diverge if not cut off and a decreasing function. Abruptly cut off power laws may be fit by log-normal functions, although not closely, and the present method may be applied.
A general form of a log-normal distribution of waiting time $\Delta t$ between events is
$f(\Delta t)=A\exp{[-(\ln{\Delta t}-\ln{\Delta t_{0}})^{2}/2\sigma^{2}]},$ (1)
where $\Delta t_{0}$ is the peak of the distribution and $\sigma$ is its dimensionless standard deviation. If there are $N$ values of $\Delta t$ in the data the normalization factor $A=N/(\sqrt{2\pi}\sigma)$.
For data with the empirical distribution $f(\ln{\Delta t})$, $\ln{\Delta t_{0}}$ is taken to be the mean
$\ln{\Delta t}_{0}={\int_{-\infty}^{\infty}\!f(\ln{\Delta t})\ln{\Delta t}\,d\ln{\Delta t}\over\int_{-\infty}^{\infty}\!f(\ln{\Delta t})\,d\ln{\Delta t}}.$ (2)
The standard deviation
$\sigma=\sqrt{{1\over\pi}{\int_{-\infty}^{\infty}\!f(\ln{\Delta t})(\ln{\Delta t}-\ln{\Delta t}_{0})^{2}\,d\ln{\Delta t}\over\int_{-\infty}^{\infty}\!f(\ln{\Delta t})\,d\ln{\Delta t}}}.$ (3)
4 Simple Models
4.1 Exact Periodicity
One limiting case of a waiting time distribution is that of periodic pulses, all of which are bright enough to be detected; all waiting times are then the same. This is often a excellent approximation for radio pulsars, whose period derivatives $10^{-21}\lesssim{\dot{P}}\lesssim 10^{-9}$. Then the fractional variation in inter-pulse intervals over a span $T$ of observations (generally, very intermittent rather than continuous) is $T{\dot{P}}/P$ and the fitted $\sigma\sim T{\dot{P}}/P$. For observed pulsars $\sigma$ is in the range $10^{-14}$–0.03 (Manchester, et al., 2005), with the largest values for young, rapidly slowing, pulsars (like the Crab) observed for decades, and the smallest values for a recycled (low-field) pulsar briefly observed. For a steadily slowing but nearly strictly periodic phenomenon like pulsar emission, $\sigma$ is not very meaningful because of its dependence on $T$.
4.2 Shot Noise
For shot noise (Poissonian statistics) with mean rate $\nu$
$f(\Delta t)=\nu\Delta t\exp{(-\nu\Delta t)}=\exp{\left[\ln{\nu\Delta t}-\exp{(\ln{\nu\Delta t})}\right]}.$ (4)
Performing the integrals in Eq. 2
$\ln{(\nu\Delta t)_{0}}=-0.577,$ (5)
and Eq. 3
$\sigma=0.723.$ (6)
4.3 Two Widely Separated $\Delta t$
Another limiting case is that of only two possible values of $\Delta t$, $\Delta t_{1}$ and $\Delta t_{2}$:
$f(\ln{\Delta t})=a\delta(\ln{\Delta t}-\ln{\Delta t}_{1})+(1-a)\delta(\ln{\Delta t}-\ln{\Delta t}_{2}).$ (7)
Then
$\ln{\Delta t}_{0}=a\ln{\Delta t}_{1}+(1-a)\ln{\Delta t}_{2}$ (8)
and
$\sigma=\sqrt{{1\over\pi}\left[a\left(\ln{({\Delta t}_{1}/{\Delta t}_{0})}\right)^{2}+(1-a)\left(\ln{({\Delta t}_{2}/{\Delta t}_{0})}\right)^{2}\right]}.$ (9)
This is the limit of a distribution with two narrow but separated peaks.
Double-peaked distributions are frequently found for the waiting times between fast radio bursts (Katz, 2019; Li et al., 2021; Zhang et al., 2021; Aggarwal et al., 2022; Hewitt et al., 2022; Niu et al., 2022; Zhang et al., 2022; Jahns et al., 2023). The peaks at shorter waiting times may be attributable to substructure within individual bursts rather than to the intervals between distinct bursts, and are not considered further.
5 Log-Normal Fit Examples
Table 1 shows the values of $\sigma$ of waiting time distributions of shot noise and of several astronomical datasets. These include very active FRB, burst storms from SGR 1806--20 and SGR 1935$+$2154 (associated with a Galactic FRB) and microglitches of the Vela pulsar.
Process or Object $N$ $\sigma$
Periodic $\infty$ 0
Shot noise $\infty$ 0.723
FRB 20121102 1652 $1.27\pm 0.02$ Li et al. (2021)
FRB 20121102 133 $1.12\pm 0.07$ Aggarwal et al. (2022)
FRB 20121102 478 $1.22\pm 0.04$ Hewitt et al. (2022)
FRB 20201124A 996 $1.04\pm 0.03$ Niu et al. (2022)
FRB 20201124A $>800$ $1.33\pm 0.03$ Zhang et al. (2022)
SGR 1806--20 111 $3.46\pm 0.23$ Hurley et al. (1994)
SGR 1806--20 262 $3.6\phantom{0}\pm 0.13$ Göğüş et al (2000)
SGR 1935$+$2154 217 $1.08\pm 0.05$ Younes et al. (2020)
Four flare stars 88–205 $0.683\pm 0.015$ Whitsett (2023)
Vela PSR $\mu$-glitches 57 $0.79\pm 0.07$ Hurley et al. (1994)
Vela PSR $\delta{\dot{\nu}}$ 57 $0.33\pm 0.03$ Hurley et al. (1994)
Table 1: Values of $\sigma$ obtained for log-normal fits to several datasets. $N$ is the number of events in each dataset. The uncertainties of $\sigma$ are their standard errors based on the number of data (where FRB intervals are evidently double-peaked the peak at longer intervals is used). Hurley et al. (1994) gave values of $\sigma$ but the other $\sigma$ (and that for Vela PSR $\delta{\dot{\nu}}$) were measured from the published figures. The FRB (see Table 6 of Hu & Huang (2022) for a review), SGR and flare star data refer to the intervals (waiting times) between bursts or flares. For FRB the data are averages over several observing epochs by the same telescope. The Vela PSR intervals are the waiting times between its microglitches and $\delta{\dot{\nu}}$ the changes in its spindown rate following these glitches. The flare star data are averages over targets 176.01, 218.01, 1224.01 and 1450.01 observed continuously by TESS (Whitsett, 2023); the uncertainty cited is the standard error of the mean $\sigma$. Niu et al. (2022) and Zhang et al. (2022) are based on the same observing campaign.
The values of $\sigma$ in the Table indicate the width of a log normal function fitted to the empirical distribution of waiting times for these datasets, and its values for a periodic signal (like a pulsar) and for shot noise. Smaller $\sigma$ indicate a narrower distribution of waiting times (the distribution is a $\delta$-function for a strictly periodic signal) and larger values indicate a wider distribution, with intervals distributed over a wide range. The values of $\sigma$ in the table depend on the exclusion of anomalously short intervals that are attributed to substructure within a single event that may be erroneously classified as distinct outbursts. The criteria for exclusion are necessarily subjective, but plausible variations change $\sigma$ by only a few hundredths. For SGR 1806-20 $\sigma$ was calculated by the cited references. The intervals range over several orders of magnitude—it has intense burst storms separated by years of quiescence, explaining the large value of $\sigma$. Other $\sigma$ were calculated from the cited published data.
6 Discussion
This paper compares the widths of log-normal fits to interval (lag) times of several different kinds of bursting phenomena, observed over several decades by different instruments. The data and analyses are necessarily heterogeneous. When the original authors provided uncertainty estimates, those are used. When they are not available, uncertainties have been estimated from the published figures, as indicated in the Table caption. The data and their analyses are described in the papers cited here.
Inspection of these figures (the reader is referred to the cited original papers) indicates that the log-normal functions fit the data well, but it is not possible to say with confidence whether deviations from the log-normal are entirely explained by the random statistics of comparatively small samples. Even if this is not the case, the log-normals explain most of the distribution of waiting times, and are a valid characterization, the the dominant terms in a moment expansion about the mean. It is not claimed that the log-normal functions describe the data exactly (aside from the statistics of small samples) but that, considered as a leading term in an expansion, they contain fundamental information about the underlying physical processes.
The virtue of $\sigma$ as a measure of the statistics of a bursting source is its independence of the observing sensitivity if the distribution of event strengths is a power law. This is simply demonstrated: If $\sigma$ depended on a detection threshold $\cal T$, then the function $\sigma({\cal T})$ would define the characteristic signal strength equal to $\cal T$ where $\sigma$ is some specific value, such as 0.723 (its value for shot noise). That would be inconsistent with a power law distribution of signal strength, a straight line on a log-log plot of number of events vs. signal strength, with no characteristic value (Kolmogorov, 1941a, b; Katz, 1986). Most of the phenomena under consideration (astronomical fast radio bursts, soft gamma repeaters and pulsar glitches) do not have obvious characteristic strengths, at least over a range several orders of magnitude wide, but there have been exceptions (Zhang et al., 2022).
Very small ($\ll 1$) values of $\sigma$ of a waiting time distribution indicate periodicity. Values $<0.723$, the value for shot noise, indicate a memory effect, like that of a noisy relaxation oscillator, that does not produce periodicity but instead has a characteristic repetition time scale, with shorter or longer waiting times less likely than for shot noise. This is often described as quasi-periodicity, and is shown in Table 1 for the post-microglitch changes of the spin-down rate of the Vela pulsar. These $\delta{\dot{\nu}}$ have a preferred scale, although the timing of the microglitches is consistent with shot noise. True periodicity is described by $\sigma\to 0$. Although this limit cannot be achieved in a finite dataset, phenomena that are called periodic, such as binary orbits or pulsar rotation, have nonzero but very small $\sigma\ll 1$.
Values of $\sigma>0.723$ indicate varying rates of activity (shot noise has the most random possible statistics if the mean or statistically expected rate of activity is unchanging). Most of the data series shown in Table 1 have $\sigma$ larger than 0.723, indicating changing levels of activity and long term memory in the mechanism. In contrast, the smaller value of $\sigma$ for the Vela $\delta{\dot{\nu}}$ indicates a characteristic scale of the phenomenon (because these data are not time intervals, it doesn’t indicate periodicity of quasi-periodicity).
Much larger values of $\sigma$, as found for SGR 1806--20, may indicate periods of much greater (“burst storms”) and of lesser activity; long intervals are likely to be followed and preceded by long intervals, and short intervals by short intervals. The fact that SGR and FRB show periods of greater and lesser activity has long been known; Zhou et al. (2022) show an extreme example for FRB 20201124A, that resumed activity after an extended quiet period (Niu, 2022). A different toy model is discussed in Sec. 4.3, in which intervals may be either long or short, but there is no memory bunching long intervals together, nor short intervals together. The fact that $\sigma$ of SGR1935$+$2154, the only SGR associated with a FRB, is much less than that of SGR 1806--20 indicates that these are fundamentally different objects; unfortunately, insufficient data are available from other SGR to explore this further.
Table 1 shows that for the repeating FRB for which sufficient waiting time data are available, $\sigma$ has approximately the same value, indicating common underlying processes. Few other repeating FRB have sufficient data to determine $\sigma$ (unfortunately, the Galactic FRB 200428, remarkably associated with a SGR, does not). In contrast, the two SGR waiting time distributions have widths whose difference far exceeds their uncertainties; perhaps SGR 1935$+$2154, the source of FRB 200428, is fundamentally different from other SGR, even the most intense of which has not been associated with a FRB (Tendulkar, Kaspi & Patel, 2016).
The fact that different observations of the same source during periods with different mean levels of activity (this applies both to FRB and to the one SGR for which two different datasets exist) are described by very similar values of $\sigma$ indicates that this is an intrinsic property of these objects. The fact that two different FRB have similar $\sigma$ suggests that repeating FRB are qualitatively similar, despite their very great differences in repetition rate and other properties; data from more repeating FRB will be required to test this hypothesis. This conclusion does not require that the values of $\sigma$ agree within their small nominal one-standard deviation uncertainties, but only that they differ by $\ll 1$. For example, the fact that values of $\sigma$ for FRB are about 2.5 less than those of SGR 1806--20 is sufficient to demonstrate that they are not produced by objects like SGR 1806--20, and the fact that they are about 0.5 greater that that of shot noise demonstrates that the FRB have memory.
The mean value of $\sigma$ for the flare stars is close to that of shot noise, as is that for Vela micro-glitches. Unlike the other processes considered, these appear to be uncorrelated random events. This is consistent with these events occuring as the result of localized processes in distinct uncoupled small regions, of the stellar photosphere for flare stars and of the neutron star interior for micro-glitches. The flare star data comprise 104, 205, 88 and 91 flares over periods of 1035 d, 1853 d, 1848 d and 1266 d, respectively, for TESS targets 176.01, 218.01, 1224.01 and 1450.01. There is no statistical evidence for cycles of activity analogous to the Solar cycle, but this may be attributable to the limited temporal extent of the data.
The universality classes of critical phenomena (Pelissetto & Vicari, 2002) suggest a path to qualitative categorization. Phenomena of different microphysical origin are sorted into universality classes on the basis of their critical exponents. Detailed knowledge of their microphysics, such as intermolecular potentials in liquid-gas critical points or the Hamiltonian of a ferromagnet, is not required to identify these classes and to establish fundamentally common dynamics among phenomena involving different physical processes. This paper suggests that the width of a log-normal distribution is analogous to the critical exponents that characterize universality classes in renormalization group theory.
Acknowledgment
I thank N. Whitsett for permission to use unpublished data and D. Eardley, J. Goodman, R. Grober, M. Maharbiz, W. Press, A. Rollett and J. Tonry for useful discussions.
Data Availability
This theoretical study did not generate any new data.